### v1

@@ Abstract
Time variation and persistence are crucial properties of volatility that are often studied separately in oil-based volatility forecasting models. Here, we propose a novel approach that allows shocks with heterogeneous persistence to vary smoothly over time, and thus model the two together. We argue that this is important because such dynamics arise naturally from the dynamic nature of shocks in oil-based commodities. We identify such dynamics from the data using localised regressions and build a model that significantly improves volatility forecasts. Such forecasting models, based on a rich persistence structure that varies smoothly over time, outperform state-of-the-art benchmark models and are particularly useful for forecasting over longer horizons.
@@ 1 Introduction
Oil-based commodities are of paramount interest to global economic prosperity because they are the most widely used source of energy. [^1] To illustrate, more than 30% of US energy consumption used petroleum-based fuels and another 30% used natural gas. Most of this energy was used in transportation, industry and electricity generation. These figures are based on reports from the Lawrence Livermore National Laboratory: https://www.llnl.gov They are constantly changing due to the constant influx of new technologies, environmental pressures and the geopolitical importance of controlling oil supplies. In recent years, large fluctuations in energy prices have been a major concern for both market participants and regulators, as price uncertainty has a significant impact on the economy (Elder and Serletis, 2010) and measuring and predicting the variability of energy prices is essential for pricing derivatives, asset allocation or risk management, but is also a key factor in understanding fluctuations in stock prices, growth rates, inflation, employment and exchange rates (Hamilton, 1983; Kilian, 2009; Kang and Yoon, 2013). While an increasing number of authors pay considerable attention to forecasting volatility (Haugom et al., 2014; Sévi, 2014; Lu et al., 2022; Zhang et al., 2019), all approaches are challenged by severe limitations imposed by model assumptions. Determining the true data-generating process of volatility dynamics becomes a challenging and open question of great priority to market participants, financial analysts and policymakers.
@@
This paper proposes a novel representation for the non-stationary volatility of oil-based commodities that allows a researcher to identify and explore its rich time-varying heterogeneous persistence structures. We aim to identify localised persistence of volatility in oil-based commodities that will be useful for modelling and forecasting purposes. Our work is closely related to the recent work of Barunik and Vacha (2023), who propose to localise the persistence structure in time series and open new avenues for modelling and forecasting. In this paper, we develop such a model and open new avenues for modelling and forecasting in the energy economics literature.
@@
Our paper contributes to the literature exploring the persistence structure of volatility through time-varying changes in unconditional volatility, persistence and the occurrence of jumps (Granger and Ding, 1996; Granger and Hyung, 2004; Bollerslev and Engle, 1993; Barunik and Vacha, 2023). We argue that energy volatility has a rich time-varying persistence structure of shocks. To exploit these stochastic properties, we use a model that can capture the time variation of the persistence structure of volatility, namely the Time-Varying Extended Wold Decomposition (TV-EWD) model of Barunik and Vacha (2023). To identify the time variation of the stochastic properties of volatility time series, the model uses the locally stationary process proposed by Dahlhaus (1996). The localisation allows to obtain the locally extended Wold decomposition, which allows a precise identification of the time-varying persistence structure. In the empirical part, we show that the time-varying persistence structure reveals significant changes in the way shocks propagate through the volatility of energy commodities.
In the empirical part, we argue that the TV-EWD model is a useful tool for understanding and forecasting the volatility of oil-based energy commodities. Specifically, we use highly liquid crude oil (CL), natural gas (NG) and regular gasoline (RBOB) futures contracts, which account for more than 75% of the trading volume in the energy commodities market. Note that crude oil is the raw material used to produce heating oil, gasoline and other petroleum-based products. First, we identify the evolution of the time-varying persistence structure of volatility. Second, we build a model based on the identified structure and use it in a forecasting exercise. Our results indicate that forecasts based on the TV-EWD model are superior to other commonly used models, especially for longer horizon forecasts. As energy commodity prices and volatility are strongly related to the geopolitical situation, global economic conditions and agents’ sentiment, it is natural to expect that the degree of persistence of shocks will vary over time. The empirical part discusses these persistence dynamics for events such as the annexation of Crimea in 2014, the Covid pandemic in 2020 and the aggression against Ukraine in February 2022.
The rest of the paper is structured as follows. Section 2 proposes a model for the time-varying persistence of oil-based commodities based on a locally stationary process, Section 3 examines the time-varying persistence of volatility in oil-based markets, and Section 4 concludes.
@@ 2.1 Locally Stationary Volatility
2.2 Time-Varying Extended Wold Decomposition
@@ 2.4 Forecasting Procedure (Model)
3 Persistence Structure of Volatility in Oil-based Commodities
The main objective of this paper is to present an improved forecasting model for oil-based commodities. In this section, we demonstrate the importance of identifying the persistence structure on the volatility time series of oil-based commodities. Namely, we use the daily realised volatilities of three major energy commodities: crude oil (CL), natural gas (NG) and regular gasoline (RBOB). The data cover the period from January 2010 to December 2022, and realised volatilities are calculated from 5-minute returns. The left column of the figure (1) shows the time series of all three commodities.
@@ 3 Persistence Structure of Volatility in Oil-based Commodities
  • Before 2014: The volatility of CL is dominated by shocks of medium persistence (16-64 days), with transitory shocks (2-8 days) of rather low importance. In contrast, NG and RBOB volatilities are dominated by transitory shocks.
  • Feb 2014 (annexation of Crimea): Significant increase in the longer-term volatility of oil and petrol, while petrol spikes only for very short periods. In December 2014, the oil price was 50% lower than in April 2014. Thus, oil volatility is mostly driven by persistent shocks of 64-128 days, which generate about 60% of the shocks in this period. Gas and petrol also show higher persistence.
  • Onset of the Covid-19 pandemic: The beginning of 2020 is characterised by the oil price spike followed by a sharp drop, leading to very high volatility and persistence. Shocks with a persistence of 64 and 128+ days generate more than 50% of the volatility. Gasoline shows similar behaviour to oil. However, natural gas does not react as strongly to the pandemic, as we observe a mixture of transient and persistent shocks (almost homogeneous structure with equal persistence).
  • Feb 2022: After the start of the aggression in February 2022, we observe the dominance of transitory shocks (2 and 4 day persistence accounts for 40%). For oil, transitory shocks also dominate in this period, but they are not as strong (2 and 4 day persistence accounts for 35%). This is a consequence of a very unstable and rapidly changing situation. After February 2022, we clearly observe an increase in overall volatility for all three commodities.
3.1 Volatility Forecasts
@@
To compare the forecasting performance of the TV-EWD model with other commonly used models for estimating the volatility of energy commodities, we use the models that best approximate persistence or long memory behaviour, such as the Heterogeneous Autoregressive Model (HAR) of Corsi (2009) and the EWD model of Ortu et al. (2020). This model also uses the extended Wold decomposition, but faces the same difficulties in capturing the time-varying nature of volatility time series. [^3] In contrast to the paper by Ortu et al. (2020), we use a shorter in-sample period and depth of decomposition. This allows the model to partially account for changes in the persistence structure of volatility. While both models capture the unconditional long-run dependence well, they assume stationarity of volatility. To assess the dynamic behaviour, we use both the simple TVP-AR(3) model and the TV-HAR model. The latter model is powerful in capturing the time-varying structure as well as in exploiting the heterogeneous horizon structure (persistence) as it allows for daily, weekly and monthly horizons. This makes the model very versatile and usually hard to beat as it is both highly effective and parsimonious.
Note that the kernel width is set to 0.3 and depth of decomposition is 7 scales which similar for all TV-EWD model settings. This means that the longest horizon considered is 128 and more days (128+). However, The TV-AR(p) model differs for prediction horizons as well as commodities. The TV-AR(p) model for CL has $p=2$ for the 1-day-ahead forecast and $p=6$ for the 5- and 22-day-ahead forecasts. For NG the TV-AR(p) model has $p=5$ for all forecasts, and for RB the TV-AR(p) model has $p=2$ for the 1-day-ahead forecast and $p=5$ for the 5- and 22-day-ahead forecasts.
@@
EWD ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ 0.962 ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$
CL TV-AR3 ${}^{{\dagger}{\dagger}}$ ${}^{{\dagger}}$ ${}^{***}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ 1.024
TV-EWD 1.009 $*{}^{*}$ ${}^{***}$ 1.008 0.982 ${}^{**}$
EWD ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$
NG TV-AR3 ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$
TV-EWD 1.003 ${}^{**}$ ${}^{***}$ 1.006 ${}^{***}$ ${}^{***}$
EWD ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$
RB TV-AR3 ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ 0.997 ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$ ${}^{{\dagger}{\dagger}{\dagger}}$
TV-EWD 1.000 $*{}^{*}$ ${}^{***}$ 1.017 1.003 ${}^{***}$

### v2

@@ Abstract
Time variation and persistence are crucial properties of volatility that are often studied separately in energy volatility forecasting models. Here, we propose a novel approach that allows shocks with heterogeneous persistence to vary smoothly over time, and thus model the two together. We argue that this is important because such dynamics arise naturally from the dynamic nature of shocks in energy commodities. We identify such dynamics from the data using localised regressions and build a model that significantly improves volatility forecasts. Such forecasting models, based on a rich persistence structure that varies smoothly over time, outperform state-of-the-art benchmark models and are particularly useful for forecasting over longer horizons.
@@ 1 Introduction
Energy commodities are of paramount interest to global economic prosperity because they are the most widely used source of energy. [^1] To illustrate, more than 30% of US energy consumption used petroleum-based fuels and another 30% used natural gas. Most of this energy was used in transportation, industry and electricity generation. These figures are based on reports from the Lawrence Livermore National Laboratory: https://www.llnl.gov They are constantly changing due to the constant influx of new technologies, environmental pressures and the geopolitical importance of controlling oil supplies. In recent years, large fluctuations in energy prices have been a major concern for both market participants and regulators, as price uncertainty has a significant impact on the economy (Elder and Serletis, 2010) and measuring and predicting the variability of energy prices is essential for pricing derivatives, asset allocation or risk management, but is also a key factor in understanding fluctuations in stock prices, growth rates, inflation, employment and exchange rates (Hamilton, 1983; Kilian, 2009; Kang and Yoon, 2013). While an increasing number of authors pay considerable attention to forecasting volatility (Haugom et al., 2014; Sévi, 2014; Lu et al., 2022; Zhang et al., 2019), all approaches are challenged by severe limitations imposed by model assumptions. Determining the true data-generating process of volatility dynamics becomes a challenging and open question of great priority to market participants, financial analysts and policymakers.
@@
This paper proposes a novel representation for the non-stationary volatility of energy commodities that allows a researcher to identify and explore its rich time-varying heterogeneous persistence structures. We aim to identify localised persistence of volatility in major energy commodities that will be useful for modelling and forecasting purposes. Our work is closely related to the recent work of Barunik and Vacha (2023), who propose to localise the persistence structure in time series and open new avenues for modelling and forecasting. In this paper, we develop such a model and open new avenues for modelling and forecasting in the energy economics literature.
@@
Our paper contributes to the literature exploring the persistence structure of volatility through time-varying changes in unconditional volatility, persistence and the occurrence of jumps (Granger and Ding, 1996; Granger and Hyung, 2004; Bollerslev and Engle, 1993; Barunik and Vacha, 2023). We argue that energy volatility has a rich time-varying persistence structure of shocks. To exploit these stochastic properties, we use a model that can capture the time variation of the persistence structure of volatility, namely the Time-Varying Extended Wold Decomposition (TV-EWD) model of Barunik and Vacha (2023). To identify the time variation of the stochastic properties of volatility time series, the model uses the locally stationary process proposed by Dahlhaus (1996). The localisation allows to obtain the locallised extended Wold decomposition, which allows a precise identification of the time-varying persistence structure. In the empirical part, we show that the time-varying persistence structure reveals significant changes in the way shocks propagate through the volatility of energy commodities.
In the empirical part, we argue that the TV-EWD model is a useful tool for understanding and forecasting the volatility of energy commodities. Specifically, we use highly liquid crude oil, natural gas and regular gasoline futures contracts, which account for more than 75% of the trading volume in the energy commodities market. Note that crude oil is the raw material used to produce heating oil, gasoline and other petroleum-based products. First, we identify the evolution of the time-varying persistence structure of volatility. Second, we build a model based on the identified structure and use it in a forecasting exercise. Our results indicate that forecasts based on the TV-EWD model are superior to other commonly used models, especially for longer horizon forecasts. As energy commodity prices and volatility are strongly related to the geopolitical situation, global economic conditions and agents’ sentiment, it is natural to expect that the degree of persistence of shocks will vary over time. The empirical part discusses these persistence dynamics for events such as the annexation of Crimea in 2014, the Covid pandemic in 2020 and the aggression against Ukraine in February 2022. In contrast to these turbulent times, we also consider the calm period of 1990’s when the volatilty was driven by very different structure of shocks.
The rest of the paper is structured as follows. Section 2 proposes a model for the time-varying persistence of energy commodities based on a locally stationary process, Section 3 examines the time-varying persistence of volatility in oil-based markets, and Section 4 concludes.
@@ 2.1 Locally stationary volatility
2.2 Time-Varying Extended Wold Decomposition (TV-EWD)
@@ 2.4 Forecasting procedure (model)
3 Persistence structure of volatility in major energy commodities
In this section, we demonstrate the importance of identifying the persistence structure in the volatility time series of energy commodities. Namely, we use the daily realised volatilities of three major energy commodities: crude oil, natural gas and regular gasoline. From the high frequency, irregularly spaced data333 obtained from Tick Data, Inc., which uses data from Globex., we extract 5-minute intraday prices of futures contracts, using the last available price in the 5-minute window. Contracts are automatically rolled over to provide continuous price records. We compute realised variance as the sum of squared intraday returns (Andersen et al., 2003) computed from log prices and annualise it to volatility as $100\times\sqrt{252\times RV_{t}}$.
It should be noted that the character of high-frequency futures data has changed dramatically over the decades, with major changes in the trading system, the introduction of near continuous trading on the CME Globex(R) electronic trading platform in December 2006, and the collection of data with milisecond time stamps in July 2011. In addition, the energy data have experienced both calm and turbulent periods where we can expect very different shocks to drive the time series. We therefore examine the performance of our model over two different periods that reflect these changes in the data.
The first period covers January 2010 to December 2022, when energy prices were influenced by several turbulent sub-periods, such as COVID-19, and when milisecond timestamp data were collected almost all day on the electronic trading platform, resulting in large volumes. In contrast, the second period, from January 1993 to December 1999, was a calm and stable period for energy data after the decades of high volatility of the Great Inflation, and high frequency data were only available during open floor trading hours with a second time stamp. Futures were only available to customers through a broker, resulting in much lower average daily volumes compared to the upcoming periods.
3.1 Energy price volatility over the period 2010-2022
During this period, futures contracts are traded on the New York Mercantile Exchange (NYMEX) and transactions are recorded with a milisecond timestamp on a 23-hour basis with a one-hour gap in trading. We redefine the day according to the electronic trading system. In addition, we exclude trades executed on US federal holidays, 24-26 December and 31 December to 2 January, due to low liquidity on these days, which could lead to an estimation bias. The data cover the period from January 2010 to December 2022. The left column of the figure (1) shows the time series for all three commodities.
@@ 3.1 Energy price volatility over the period 2010-2022
  • Before 2014: The volatility of CL is dominated by components calculated from shocks of medium persistence (16-64 days), with transitory shocks (2-8 days) of rather low importance. In contrast, NG and RBOB volatilities are dominated by transitory shocks.
  • Feb 2014 (annexation of Crimea): The importance of the more persistent components of oil and gasoline has increased significantly, with gasoline spiking only for very short periods. In December 2014, the oil price was 50% lower than in April 2014, and oil volatility is mostly driven by persistent shocks of 64-128 days, which generate about 60% of the shocks in this period. Gas and petrol also show higher persistence.
  • Onset of the Covid-19 pandemic: The beginning of 2020 is characterised by the oil price spike followed by a sharp drop, leading to very high volatility. Interestingly, components driven by shocks with a persistence of 64 and 128+ days generate more than 50% of the volatility. Gasoline shows similar behaviour to oil. However, natural gas does not react as strongly to the pandemic, as we observe a mixture of transient and persistent shocks (almost homogeneous structure with equal persistence).
  • Feb 2022: After the start of the aggression in February 2022, we observe a strong dynamic with the dominance of persistent shocks being overtaken by transitory shocks (2 and 4 day persistence accounts for 40%). For oil, transitory shocks also dominate in this period, but they are not as strong (2 and 4 day persistence accounts for 35%). This is a consequence of a very unstable and rapidly changing situation, in which transitory shocks to market participants’ expectations generated a large proportion of price fluctuations. After February 2022, we clearly observe an increase in overall volatility for all three commodities with mixed persistence structures.
Note that although the level of volatility in these periods is very different, especially Covid-19 is characterised by very high volatility, all turbulent sub-periods are driven by shocks with higher persistence due to the increased uncertainty of market participants. Therefore, compared to other calmer periods, we can observe a higher proportion of higher persistence components driving the series. We will return to this discussion in a later section that examines the calm period of the 1990s. It is also interesting to note that this is not the case for natural gas, which is driven by transitory components after February 2022. With the immense impact of the Russian invasion on the importance of gas, we observe large fluctuations in volatility after February 2022, showing that a lot of transitory uncertainty is driving the time series.
3.1.1 Volatility forecasts
@@
To compare the forecasting performance of the TV-EWD model with other commonly used models for estimating the volatility of energy commodities, we use the models that best approximate persistence or long memory behaviour, such as the Heterogeneous Autoregressive Model (HAR) of Corsi (2009) and the EWD model of Ortu et al. (2020). This model also uses the extended Wold decomposition, but faces the same difficulties in capturing the time-varying nature of volatility time series. [^4] In contrast to the paper by Ortu et al. (2020), we use a shorter in-sample period and depth of decomposition. This allows the model to partially account for changes in the persistence structure of volatility. While both models capture the unconditional long-run dependence well, they assume stationarity of volatility. To assess the dynamic behaviour, we use both the simple TVP-AR(3) model and the TV-HAR model. The latter model is powerful in capturing the time-varying structure as well as in exploiting the heterogeneous horizon structure (persistence) as it allows for daily, weekly and monthly horizons. This makes the model very versatile and usually hard to beat as it is both highly effective and parsimonious. We save the first 700 observations for the in-sample fit and obtain 2600 forecasts from a rolling window starting the out-of-sample forecasting period with September 15, 2012.
Note that the kernel width is set to 0.3 and the decomposition depth to 7 scales, which is similar for all TV-EWD model settings. This means that the longest horizon considered is 128 days and more (128+). However, the TV-AR(p) model differs for both forecast horizons and commodities. The TV-AR(p) model for CL has $p=2$ for the 1-day-ahead forecast and $p=6$ for the 5- and 22-day-ahead forecasts. For NG the TV-AR(p) model has $p=5$ for all forecasts and for RB the TV-AR(p) model has $p=2$ for the 1-day-ahead forecast and $p=5$ for the 5- and 22-day-ahead forecasts. All parameters, such as the number of lags of the approximating model, the kernel widths or the degree of decomposition, have been chosen with a view to improving the out-of-sample performance for a given time series and thus reducing losses. One of the key parameters of the model is the number of lags $p$ in the TV-AR($p$), which determines the decomposition. For longer horizon forecasts, it is crucial to choose a larger $p$ as it helps to capture a richer structure in terms of persistence. On the other hand, for $h=1$ forecasts, a more parsimonious model with fewer lags helps to reduce the error. Again, this parameter is selected for each forecast horizon and time series based on the largest reduction in losses.
@@
EWD 1.373††† 1.542††† 0.962 1.457††† 1.577††† 1.185†††
CL TV-AR3 1.051†† 1.062 0.855∗∗∗ 1.035††† 1.108††† 1.024
TV-EWD 1.009 0.946 0.807∗∗∗ 1.008 0.982 0.960∗∗
EWD 1.297††† 1.381††† 1.235††† 1.437††† 1.621††† 1.358†††
NG TV-AR3 1.064††† 1.134††† 1.300†† 1.089††† 1.253††† 1.234†††
TV-EWD 1.003 0.945∗∗ 0.923∗∗∗ 1.006 0.965∗∗∗ 0.895∗∗∗
EWD 1.249††† 1.444††† 1.125††† 1.481††† 1.535††† 1.299†††
RB TV-AR3 1.052††† 1.112††† 0.997 1.084††† 1.186††† 1.105†††
TV-EWD 1.000 0.954 0.880∗∗∗ 1.017 1.003 0.951∗∗∗
@@
3.2 Energy price volatility of the 1993-1999 period
While 2010-2022 period was characterised by more persistent shocks driving energy volatility, together with turbulent times and high volumes, we would like to challenge our model with the calm period of the 1990s, when energy futures enjoyed much less volatile prices and much lower volumes. The trades were recorded with a second time stamp during a regular hour when the floor was open. Note that we use unleaded gasoline (HU) futures instead of RBOB gasoline futures, which were introduced after December 2006 only. During this period, we also exclude trades executed on US federal holidays, 24-26 December and 31 December to 2 January, due to low liquidity on these days, which could lead to an estimation bias. Data cover the period from January 1993 to December 1999. The left column of the figure (2) shows the time series for all three commodities, and it is immediately apparent that the series are much more stable and less volatile.
Figure 2: Left: Realized volatilities for crude oil (CL), natural gas (NG) and gasoline (HU). Right: Time-varying persistence structure of CL (top), NG (middle), HU (bottom). The plot shows the ratios of $\widehat{\beta}^{j}(t/T,1)/\sum_{j}\widehat{\beta}^{j}(t/T,1)$ on the y-axis, with $j$ corresponding to 2, 4, 8, 16, 32, 64, 128+ days persistence of the shocks represented by cold to warm colours, respectively, over the period from January 1993 to December 1999 on the x-axis.
Similarly to the previous period, we use the multiscale impulse response functions from the TV-EWD model to capture the evolution of the persistence structure of volatility and the persistence of the shocks that generate the volatility time series over this period. Similarly to the previous section, we report the persistence at a given scale $j$ as the persistence share of that scale relative to the persistence at other scales.
Figure (2) shows a stark contrast in the persistence structures compared to the previous turbulent period 2010-2022 discussed in the previous section. The main change is that components with higher persistence play a weaker role in the time series and, conversely, the series seem to be driven by transitory shocks. This is important information for modelling purposes.
3.2.1 Volatility forecasts
The quieter period 1993-1999 is mainly driven by transitory shocks, but at the same time we can observe some change in the structure. Therefore, our TV-EWD model, which is able to capture the non-trivial dynamics, can also be useful for forecasting, although with increasing complexity it may lose its forecasting power, especially because the model is useful for data with strong long-term persistence.
To compare the forecasting performance of the TV-EWD model, we again use the same models in the same setting, saving the first 700 observations for the in-sample fit and leaving the rest for the out-of-sample forecast period.
Note that the kernel width is set to 0.3 and the depth of decomposition to 7 scales, which is similar for all TV-EWD model settings. This means that the longest horizon considered is 128 and more days (128+). However, the TV-AR(p) model differs for both forecast horizons and commodities. The TV-AR(p) model for CL has $p=3$ for all forecasts. For NG, the TV-AR(p) model has $p=5$ for the 1- and 5-day-ahead forecasts and $p=3$ for the 22-day-ahead forecast, and for HU the TV-AR(p) model has $p=2$ for the 1-day-ahead forecast and $p=5$ for the 5- and 22-day-ahead forecasts.
Table 2 summarises the results of volatility forecasting performance over this period. Again, the TV-HAR model is used as the benchmark and all reported losses are relative to the TV-HAR. In general, the TV-EWD model’s ability to outperform all alternatives is weaker during this calm period compared to the turbulent period of 2010-2022. We attribute this mainly to the fact that the high persistence component is weak and therefore time series driven by transitory shocks are more difficult to forecast over multiple horizons. TV-EWD provides improved forecasts in all cases, but the forecast errors are not always significantly different from other models.
[…]
RMSE MAE
$h=1$ $h=5$ $h=22$ $h=1$ $h=5$ $h=22$
EWD 1.122††† 1.206††† 1.231††† 1.128††† 1.301††† 1.354†††
CL TV-AR3 1.013 1.029 0.972 1.024†† 1.062††† 1.013
TV-EWD 0.995 0.975 0.978 0.990 0.998 1.002
EWD 1.055††† 1.005 0.946∗∗ 1.093††† 1.056†† 0.914∗∗∗
NG TV-AR3 1.032††† 1.081††† 1.065††† 1.050††† 1.158††† 1.063†††
TV-EWD 1.002 0.968 0.937∗∗∗ 1.003 0.995 0.960∗∗
EWD 1.080††† 1.128††† 1.247††† 1.106††† 1.181††† 1.351†††
HU TV-AR3 1.014 1.042 1.047 1.025†† 1.093††† 1.134†††
TV-EWD 0.998 0.964 0.918 0.998 0.979 0.976